Two identical small current carrying loops of wire
Answer #1

a) By definition: \mu =IA=I\cdot \pi R^{2} (it is a vector, but here is the formula for its modulus)

b) B=\frac{\mu _{0}R^{2}I}{2\left ( x^{2} +R^{2}\right )^{3/2}} (there is a deduction; ife necessary, I'll give it here)

\mu _{0}=permitivity of vacuum

c) Use the formula for F given in the text, and get:

\left | F \right |=\frac{3\pi \mu _{0}R^{4}I}{2}\cdot \frac{x\sqrt{x^{2}+R^{2}}}{\left ( x^{2} +R^{2}\right )^{3}}\: \left ( 1 \right )

d) Extract x2 from the radical at numerator, and x2 - forced factor from denominator. We'll get:

\left | F \right |=\frac{3\pi \mu _{0}R^{4}I}{2}\cdot \frac{x^{2}\sqrt{1+\frac{R^{2}}{x^{2}} }}{\left [x^{2} \left ( 1+\frac{R^{2}}{x^{2}} \right )\right ]^{3}}\sim \frac{1}{x^{4}}

[the radical from the numerator=1; the small bracket from the denominator=1]

d) Replace the numbers in (1) and get: F~2.3914*10-13 N.